Session: Nov 1997 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Venturi Tiziana at Wed Nov 26 00:09:28 MET 1997) Minor failures
VC02 has no power, due to the fact that the frequency was outside the range of the receiver. Other than that the experiment ended without apparent problems.

Torun (K.Borkowski at Wed Nov 26 09:24:57 MET 1997) Minor failures
The observations went smoothly, but the observers, curiously enough, used a tape that was previously partially recorded during the terminal tests. Thus sections of these observations may be of poorer quality.

Westerbork (tony foley at Wed Nov 26 11:08:13 MET 1997) Success
No major problems but some strong EMI in the VC1 and VC2, particularly around the end of the run (2245UT)

Noto (C.Trigilio at Wed Nov 26 12:38:17 MET 1997) Success

Onsala (John Conway at Wed Nov 26 18:58:47 MET 1997) Minor failures
No known problems. However since we only have 4 external 250kHz filters available 2 channels were not observed. The 4 filters were placed one on each line at RCP plus one at 1612 LCP.

Effelsberg (dgraham at Thu Nov 27 09:48:28 MET 1997) Success

Jodrell (Alastair Gunn at Thu Nov 27 12:43:59 MET 1997) Minor failures
Some RFI. Tape stopped several times during this experiment, reason unknown. Tape came off spool and schedule restarted at 17:34. Some cable problem may have affected attenuation on IF2.

Calibration (Fredrik T Rantakyro at Tue Dec 9 15:02:13 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name gs013.antab Read the comments in the file to get further information.

MPIfR (Walter Alef at Wed Jan 7 11:48:48 MET 1998) Failure
Correlation cancelled due to MK4 formatter bug: 500 kbit sampling rate does not work.